# Text2SQL-Flow: A Robust SQL-Aware Data Augmentation Framework for Text-to-SQL

### v2

@@ III-A1 SQL Augmentation
  • Data Value Transformations: Adjust filtering conditions, date ranges, or numeric thresholds to refine query precision; modify sorting criteria or limit values; alter aggregation boundaries (e.g., GROUP BY with different temporal granularities).
  • Query Structure Modifications: Rewrite aggregation queries into window functions and vice versa; restructure simple queries into subqueries or common table expressions (CTEs); substitute JOIN with EXISTS or IN; switch between correlated and uncorrelated subqueries to enhance flexibility and maintainability.
  • Business Logic Changes: Adapt queries to alternative business domains, such as shifting from sales analysis to inventory management or from customer analytics to supplier evaluation; adjust data granularity (e.g., daily data to monthly summaries); change the analysis perspective (e.g., profit analysis to cost analysis); or modify metrics (e.g., sum to average, count to percentage).
  • Complexity Enhancements: Introduce additional filtering conditions, join extra tables, incorporate CASE expressions for conditional logic, or embed data validation and quality checks to increase query complexity.
  • Advanced SQL Features: Utilize partitioned window functions, UNION/INTERSECT/EXCEPT operations, recursive CTEs, and pivot/unpivot operations to broaden the technical scope of queries.
  • Performance and Optimization: Improve execution efficiency by adding optimization hints, restructuring queries to leverage indexes, adopting efficient query patterns, and optimizing WHERE clauses.
To ensure the validity of the generated augmented SQL queries, we incorporate randomly sampled database values into the prompts to improve semantic grounding. Each prompt consists of the following components: (1) Instruction ($I_{\text{sql}}$): a directive guiding the LLM to generate SQL queries that satisfy realistic analytical requirements; (2) Database Schema ($S$): the full set of CREATE TABLE statements describing all relations in the database; (3) Database Values ($V$): a set of randomly sampled (column, value) pairs drawn from the underlying tables to instantiate meaningful query predicates; (4) Original SQL query ($s_{i}^{\text{orig}}$): the query to be augmented; and (5) Augmentation Direction ($c$): one of six predefined augmentation strategies. Let $C=\{c_{1},\dots,c_{6}\}$ denote the set of augmentation strategies. For each original query $s_{i}^{\text{orig}}$, we uniformly sample a strategy $c\sim\text{Uniform}(C)$, draw representative database values $V$ from the relevant tables, and construct an augmentation prompt $P_{\text{aug}}(I,S,s_{i}^{\text{orig}},V,c)$. The augmented query is then produced by the language model as:
@@
where $P_{\text{aug}}(\cdot)$ denotes the deterministic prompt construction process, and $\text{LLM}(\cdot)$ represents the stochastic query generation by the language model.
@@
III-A2 SQL Execution Filter
Not all generated SQL queries are valid or efficient. Therefore, it is necessary to filter the SQL queries. An SQL execution filter is applied to screen out unsuitable candidates. The filter evaluates queries along two dimensions: (1) Executability, which verifies that the SQL syntax is correct and the query can be successfully executed on the target database; and (2) Execution Efficiency, which discards queries whose runtime exceeds a predefined threshold to ensure system responsiveness. After applying this filter, all resulting SQL queries are guaranteed to be valid.
III-A3 Question Generation
After obtaining the valid augmented SQL queries, the next step is to generate semantically equivalent NL questions. To ensure linguistic diversity in the synthetic data, it is essential to incorporate a wide range of stylistic instructions. Inspired by OmniSQL [li2025omnisql], we analyze real-world user queries along multiple linguistic dimensions and identify eleven common stylistic categories, grouped into four high-level aspects:
  • Tone and Formality: formal and colloquial
  • Syntactic Structure and Intent: imperative, interrogative, and declarative
  • Information Density and Clarity: concise, descriptive, vague, and metaphorical
  • Interaction Patterns: role-playing and procedural
@@ III-A3 Question Generation
To ensure both semantic fidelity and linguistic diversity, we generate $k$ candidate questions $\{q_{i1},\dots,q_{ik}\}$ for each $s_{i}^{\text{aug}}$ by independently sampling a new style $\tau$ for each generation. Among these candidates, we compute pairwise semantic similarities and retain the question with the highest average similarity to the others, thereby preserving semantic consistency while encouraging stylistic variation. The final synthetic dataset is constructed as $\mathcal{D}=\{(q_{i},s_{i}^{\text{aug}})\}_{i=1}^{N}$, where $N$ is the number of distinct SQL queries. This approach yields a dataset with rich linguistic variation and broad semantic coverage, enhancing both robustness and generalization in downstream applications.
III-A4 Chain-of-Thought Generation
Chain-of-Thought (CoT) reasoning enhances a model’s ability to solve complex tasks by decomposing them into a sequence of smaller, manageable subproblems, thereby enabling step-by-step reasoning. Equipping LLMs with CoT reasoning capabilities thus improves their overall reasoning performance. In the context of Text-to-SQL translation, the CoT reasoning process typically involves the following steps: (1) analyzing the intent of the NL question, (2) interpreting the database schema, (3) identifying relevant tables and columns, (4) filtering necessary information, (5) selecting appropriate SQL operators, and (6) incrementally constructing the final SQL.
To generate CoT reasoning traces, we employ an LLM with strong reasoning capabilities. The prompt design typically includes (1) instruction($I_{\text{cot}}$), (2) database schema($S$), (3) generated NL question($q_{i}$), and (4) augmented SQL query($s_{i}^{\text{aug}}$). Guided by this prompt for CoT generation, the LLM produces a complete reasoning chain that encompasses intermediate reasoning steps as well as the final SQL query, which can be expressed as follows:
$cot_{i}=\text{LLM}\big(P_{\text{cot}}(I_{\text{cot}},S,s_{i}^{\text{aug}},q_{i})\big).$ (3)
where $P_{\text{cot}}(\cdot)$ denotes the process of CoT generation. During validation of the CoT, the generated SQL query is extracted from the reasoning chain. The CoT process is accepted as a valid solution only if the execution result of the generated SQL matches that of the reference SQL on the given database.
III-A5 Prompt Generation
The prompt ($p_{i}$) serves as the primary input to the model, containing all necessary information for it to perform reasoning. To facilitate reliable text-to-SQL generation, a well-structured prompt incorporates not only the NL question ($q_{i}$) but also the database schema ($S$) and a specific task instruction ($I_{\text{prop}}$) that guides the model. The composition process is formally defined as:
@@ III-A5 Prompt Generation
where $P_{\text{prop}}(\cdot)$ denotes the function that synthesizes the final prompt from its constituent parts.
III-A6 SQL Component Classifier
$cd_{i}=\text{CC}(s_{i}^{\text{aug}})$ (5)
where $\text{CC}(\cdot)$ denotes the mapping function of the Component Classifier, and $cd_{i}$ represents the component difficulty of the query.
III-A7 SQL Execution Classifier
$ed_{i}=1-\frac{n}{k}$ (6)
Figure 4: Comparison of visualization between original SQLs and augmented SQLs using our framework.
III-A8 Augmented Data Overview
@@
Leveraging the proposed Text-to-SQL data augmentation pipeline, Text2SQL-Flow, we generate an augmented dataset from the original SQL-query pairs $s_{i}^{\text{ori}}$. The pipeline systematically synthesizes diverse and valid Text-to-SQL examples, which can serve as a valuable supplement to existing Text-to-SQL datasets. The final dataset comprises $N$ entries of the following form:
@@
Using Text2SQL-Flow, we construct a dataset named SQLFlow. The raw seed data is sourced from widely used benchmarks. The final dataset comprises a total of 89,544 entries: 37,517 derived from the Spider-train dataset [yu2018spider], 37,536 from the BIRD-train dataset [li2024can], and 14,491 from the EHRSQL-train dataset [lee2022ehrsql]. In this paper, we refer to these subsets as SQLFlow-Spider, SQLFlow-BIRD, and SQLFlow-EHRSQL, respectively. It should be noted that these data are augmented and do not overlap with the original training datasets. Our approach augments existing public datasets, thereby enriching their diversity and coverage. It is scalable, efficient, and capable of generating high-quality synthetic data. As shown in Table I, our pipeline not only mitigates the sparsity and structural gaps inherent in the original dataset but also substantially extends its complexity frontier—surpassing the original in both scale and quality. These improvements are visually illustrated in Figure 4, which presents a 2D t-SNE scatter plot generated from feature vectors derived from the data summarized in Table I.
III-B Database Manager System Design
Part III-A describes the data augmentation pipeline; however, within the framework, an efficient and reliable database interaction mechanism serves as the core infrastructure underpinning the stable operation of the entire pipeline. Specifically, multiple critical steps within the pipeline, such as SQL execution filtering and SQL augmentation, require access to the underlying database system. These interactions can be categorized into two types: (1) SQL Execution and (2) Schema Metadata Extraction.
SQL Execution involves submitting generated SQL queries to the database engine and retrieving the execution feedback, which is primarily used to validate the actual executability of the SQL queries. However, mainstream database systems, such as SQLite, MySQL, and PostgreSQL, exhibit significant differences in their driver interfaces and connection protocols. Directly invoking each database’s native API not only complicates cross-platform compatibility but also necessitates maintaining separate adaptation logic for every system, leading to code redundancy and severely limiting the system’s scalability.
Schema Metadata Extraction aims to retrieve structured metadata about the database, including table definitions like CREATE TABLE statements, column attributes (data type, nullability, default values), primary/foreign key constraints, index information, and representative INSERT statements that can be used for SQL augmentation. Such metadata is foundational for understanding database semantics, generating syntactically and semantically valid SQL queries, and implementing context-sensitive data augmentation strategies. Although most relational databases provide mechanisms, such as querying INFORMATION_SCHEMA or system catalog tables, to retrieve schema information, the query syntax, field naming conventions, and data formats vary significantly across systems, further complicating unified handling. Moreover, for a given database instance, the schema typically remains static over short time intervals; repeatedly issuing identical metadata queries introduces unnecessary I/O overhead and latency, substantially degrading the overall throughput of the pipeline.
To systematically address these challenges, we designed and implemented a Database Manager module within the underlying architecture of the data augmentation framework. This module abstracts away low-level database interaction details and provides upper-layer components with a unified, efficient, and extensible programming interface. Specifically, the Database Manager supports batch_sql_execution and batch_compare_sql, improving processing throughput under high-concurrency scenarios. Additionally, it encapsulates schema metadata retrieval logic and exposes high-level methods such as get_ddl (to obtain Data Definition Language) and get_insert_statement (to generate representative insertion statements), thereby simplifying upper-layer dependencies on database structure.
To achieve true cross-database compatibility, we introduced an abstract base class DatabaseConnector. This class explicitly defines a standardized interface, including methods such as connect_db (establish connection and perform initialization), execute_sql (execute arbitrary SQL and return results), and get_schema (extract complete schema metadata). For each specific database system, one only needs to inherit from this base class and implement its database-specific driver calls and error-handling logic (e.g., SQLiteDatabaseConnector). This enables seamless integration into the system. At runtime, the Database Manager dynamically instantiates the appropriate connector based on configuration, while its core logic remains agnostic to the underlying database type, thus achieving the architectural goal of “design once, adapt to many databases.”
To further enhance overall throughput, the Database Manager incorporates a cache-based schema management mechanism. Upon the first request for a database’s schema, the system serializes its structured metadata and stores it in an in-memory cache. Subsequent requests for the same schema are served directly from the cache, eliminating redundant queries and reducing I/O load and response latency. This mechanism is especially effective in large-scale SQL generation and validation tasks, where hundreds to thousands of repeated schema queries against the same database are common.
III-C Synthetic Data for Open-Source Models
For open-source LLMs, their reasoning capabilities can be enhanced through supervised fine-tuning (SFT), a process in which the model learns to map inputs to outputs. In the case of augmented data from Text2SQL-Flow, each instance includes a prompt $p_{i}$ and a corresponding CoT $cot_{i}$. The pairs $\langle p_{i},cot_{i}\rangle$ extracted from the SQLFlow dataset serve as high-quality training data $\mathcal{D}$. During SFT, we train on the full output sequence using standard causal language modeling:
@@ III-C Synthetic Data for Open-Source Models
where the input $x$ corresponds to the prompt $p_{i}$, and the output $y$ represents the complete CoT sequence $cot_{i}$. The model parameters are denoted by $\theta$. This loss function, applied over the entire output sequence, encourages the model not only to generate correct SQL queries but also to internalize interpretable and structured reasoning pathways.
III-D Synthetic Data for Prompting-based Methods
@@
However, existing methods for retrieving few-shot examples fail to provide effective guidance because similarity has not been sufficiently exploited. In this case, we propose a structure-aware few-shot retrieval method for Text-to-SQL. The core idea is to map both NL questions and SQL queries into a unified, detail-abstracted semantic space, where retrieval is performed based on structural similarity. Specifically, inspired by the masking strategy in DAIL-SQL [gao2023text], we apply standardized masking to schema-dependent elements (e.g., table/column names), string literals, and numeric constants in both questions and SQL queries, denoted as $\text{Mask}(q)$ and $\text{Mask}(\text{s})$, respectively. Building upon this, we develop a structure-aware embedding model based on an LLM to compute semantic similarity between masked NL questions and masked SQL queries. Given a target question $q_{t}$, we retrieve the top-$k$ examples from a knowledge base $B=\{(q_{i},s_{i})\}_{i=1}^{N}$ whose masked SQL representations $\text{Mask}(s_{i})$ are closest to $\text{Mask}(q_{t})$ in the embedding space, and use the original $(q_{i},s_{i})$ pairs as the final few-shot prompts. The knowledge base $B$ is generated based on the SQLFlow dataset we construct.
@@
In our implementation, we adopt the Qwen3-Embedding architecture [zhang2025qwen3], employing a causal-attention-based LLM as the embedding encoder. Both the knowledge base and the training data for the embedding model are derived from the SQLFlow dataset we constructed. The input format concatenates the task instruction with the masked question as follows: $\texttt{\{Instruction\} \{Mask}(q_{i})\texttt{\}}$ followed by an [EOS] token. The final sentence embedding is obtained from the hidden state of the [EOS] token in the last layer of the encoder. Our method retrieves few-shot examples that are highly aligned with the target query in SQL structure. This enhances the few-shot reasoning capability of closed-source LLMs on the Text-to-SQL task.

### v3

@@ III-B1 SQL Augmentation
  • Data Value Transformations (DVT): Modify query parameters to refine result precision, including filtering conditions, date ranges, numeric thresholds, sorting criteria, limit values, and aggregation boundaries (e.g., GROUP BY at different temporal granularities).
  • Query Structure Modifications (QSM): Alter the structural form of queries by rewriting aggregation queries as window functions (and vice versa), introducing subqueries or common table expressions (CTEs), substituting JOIN with EXISTS or IN, and switching between correlated and uncorrelated subqueries to improve flexibility and maintainability.
  • Business Logic Changes (BLC): Adapt queries to alternative analytical objectives or business contexts, such as shifting across domains (e.g., sales to inventory), adjusting data granularity (e.g., daily to monthly), changing analytical perspectives (e.g., profit to cost), or modifying evaluation metrics (e.g., sum to average).
  • Complexity Enhancements (CE): Increase query complexity by introducing additional filtering conditions, joining extra tables, incorporating CASE expressions for conditional logic, or embedding data validation and quality checks.
  • Advanced SQL Features (ASF): Employ advanced SQL constructs, including partitioned window functions, set operations (UNION, INTERSECT, EXCEPT), recursive CTEs, and pivot/unpivot operations, to extend the technical expressiveness of queries.
  • Performance and Optimization (PO): Optimize query execution by applying optimization hints, restructuring queries to leverage indexes, adopting efficient query patterns, and refining WHERE clauses to balance performance and resource utilization.
To ensure the validity of the generated augmented SQL queries, we incorporate randomly sampled database values into the prompts to improve semantic grounding. Each prompt consists of the following components: (1) Instruction ($I_{\text{sql}}$): a directive guiding the LLM to generate SQL queries that satisfy realistic analytical requirements; (2) Database Schema ($S$): the full set of CREATE TABLE statements describing all relations in the database; (3) Database Values ($V$): Values are randomly sampled from multiple rows in each table to construct meaningful query predicates; any sampled row containing null values is replaced. (4) Original SQL query ($s_{i}^{\text{orig}}$): the query to be augmented; and (5) Augmentation Direction ($c$): one of six predefined augmentation strategies. Let $C=\{c_{1},\dots,c_{6}\}$ denote the set of augmentation strategies. For each original query $s_{i}^{\text{orig}}$, we uniformly sample a strategy $c\sim\text{Uniform}(C)$, draw representative database values $V$ from the relevant tables, and construct an augmentation prompt $P_{\text{aug}}(I,S,s_{i}^{\text{orig}},V,c)$. The augmented query is then produced by the LLM as:
@@
where $P_{\text{aug}}(\cdot)$ denotes the prompt construction process, and $\text{LLM}(\cdot)$ represents the stochastic query generation by LLM.
@@
III-B2 SQL Executability Filter
Not all generated SQL queries are executable, as some may contain syntax errors or reference non-existent database objects. Therefore, we apply an SQL executability filter to remove invalid candidates. Under a fixed database schema and a read-only setting without user-defined functions, we verify executability at the plan level. Specifically, for each generated SQL query, we check whether the database can successfully parse the query and generate an execution plan using EXPLAIN (e.g., EXPLAIN QUERY PLAN in SQLite). A query is retained if plan generation succeeds; otherwise, it is discarded. Since this verification does not execute the query, it is insensitive to runtime variability such as cold caches or transient system load. This design choice avoids unintentionally biasing the augmented SQL set toward “fast” queries, while ensuring that all retained queries are logically executable under the fixed schema, which is sufficient for our SQL data augmentation task.
III-B3 Question Generation
After obtaining the valid augmented SQL queries, the next step is to generate semantically equivalent NL questions. To ensure linguistic diversity in the synthetic data, it is essential to incorporate a wide range of stylistic instructions. Inspired by OmniSQL [29], we analyze real-world user queries along multiple linguistic dimensions and identify eleven common stylistic categories, grouped into four high-level aspects:
  • Tone and Formality (TF): levels of formality, including formal and colloquial styles
  • Syntactic Structure and Intent (SSI): sentence types such as imperative, interrogative, and declarative
  • Information Density and Clarity (IDC): degrees of conciseness, descriptiveness, vagueness, and metaphorical expression
  • Interaction Patterns (IP): interaction modes such as role-playing and procedural interactions
@@ III-B3 Question Generation
To ensure both semantic fidelity and linguistic diversity, we generate $k$ candidate questions $\{q_{i1},\dots,q_{ik}\}$ for each $s_{i}^{\text{aug}}$ by independently sampling a new style $\tau$ for each generation. Among these candidates, we select the candidate question that maximizes semantic consistency with the underlying SQL query, while maintaining agreement across stylistic variants. The final synthetic dataset is constructed as $\mathcal{D}=\{(q_{i},s_{i}^{\text{aug}})\}_{i=1}^{N}$, where $N$ is the number of distinct SQL queries. This approach yields a dataset with rich linguistic variation and broad semantic coverage, enhancing both robustness and generalization in downstream applications.
III-B4 NL–SQL Correspondence Filter
Even when an SQL query is executable, it may not faithfully reflect the intent of its paired NL question. To ensure semantic consistency between the NL question and the corresponding SQL query, we introduce a NL–SQL correspondence filter. Specifically, given a database schema $S$, a natural language question $q_{i}$, and its augmented SQL query $s_{i}^{\text{aug}}$, we employ LLM to assess whether the two express the same query intent:
$\mathbb{I}_{\text{align}}(S,s_{i}^{\text{aug}},q_{i})=\mathbb{I}_{\text{LLM}}(S,s_{i}^{\text{aug}},q_{i}),$ (4)
where $\mathbb{I}_{\text{LLM}}(\cdot)\in\{0,1\}$ indicates whether semantic equivalence is confirmed. Only question–SQL pairs with $\mathbb{I}_{\text{align}}=1$ are retained for subsequent use.
III-B5 Prompt Generation
For open-source LLMs, problem-solving capabilities can be enhanced through supervised fine-tuning (SFT), a process in which the model learns to map prompt inputs $p_{i}$ to the corresponding outputs. To facilitate reliable Text-to-SQL generation, a well-structured prompt ($p_{i}$) incorporates not only the NL question ($q_{i}$) but also the database schema ($S$) and a specific task instruction ($I_{\text{prop}}$) that guides the model. The process of prompt construction is formally defined as follows:
@@ III-B5 Prompt Generation
where $P_{\text{prop}}(\cdot)$ denotes the function that synthesizes the final prompt from its constituent components.
III-B6 Chain-of-Thought Generation
To train high-quality LLMs capable of solving complex problems, supervising models using only the final SQL outputs is often insufficient. Chain-of-thought (CoT) traces generated by stronger models decompose the question-solving process into a sequence of smaller, manageable subproblems, enabling step-by-step reasoning. Such CoT supervision provides informative learning signals and can be leveraged to improve the overall reasoning performance of weaker LLMs. In the context of Text-to-SQL translation, CoT traces typically involve the following steps: (1) analyzing the intent of the NL question, (2) interpreting the database schema, (3) identifying relevant tables and columns, (4) filtering necessary information, (5) selecting appropriate SQL operators, and (6) incrementally constructing the SQL.
To generate CoT reasoning traces, we employ an LLM with strong reasoning capabilities. The prompt design includes (1) an instruction ($I_{\text{cot}}$), (2) the database schema ($S$), (3) the generated NL question ($q_{i}$), and (4) the augmented SQL query ($s_{i}^{\text{aug}}$). Guided by this prompt, the LLM produces a complete reasoning chain that encompasses intermediate reasoning steps as well as the final SQL query, which can be expressed as:
$cot_{i}=\text{LLM}\big(P_{\text{cot}}(I_{\text{cot}},S,s_{i}^{\text{aug}},q_{i})\big)$ (6)
where $P_{\text{cot}}(\cdot)$ denotes the CoT prompt construction process. During validation, the generated SQL query is extracted from the reasoning chain. A CoT trace is accepted as a valid solution only if the execution result of the generated SQL matches that of the reference SQL on the given database.
Figure 4: Visualization of changes in distribution-level (left) and query-level (right) diversity induced by data augmentation.
III-B7 Augmented Data Overview
@@
Leveraging the proposed Text-to-SQL data augmentation pipeline, we generate an augmented dataset from the original SQL-query pairs $s_{i}^{\text{ori}}$. The pipeline systematically synthesizes diverse and valid Text-to-SQL examples, which can serve as a valuable supplement to existing Text-to-SQL datasets. The final dataset comprises $N$ entries of the following form:
$\left\{s_{i}^{\text{aug}},\,q_{i},\,S,\,p_{i},\,\text{cot}_{i}\right\}_{i=1}^{N}$
@@
Using Text2SQL-Flow, we construct a dataset named SQLFlow. The raw seed data is sourced from widely used benchmarks. The final dataset comprises a total of 75,386 entries: 30,502 derived from the Spider-train dataset [52], 33,163 from the BIRD-train dataset [32], and 11,721 from the EHRSQL-train dataset [26]. In this paper, we refer to these subsets as SQLFlow-Spider, SQLFlow-BIRD, and SQLFlow-EHRSQL, respectively. It should be noted that the augmented data do not overlap with the original training datasets. In addition, an n-gram-based filtering strategy is applied to ensure that SQLFlow has no overlap with the test data, thereby preventing any potential data leakage. Overall, our approach augments existing public datasets and is scalable, efficient, and capable of generating high-quality synthetic data. As shown in Table I, our pipeline not only mitigates the sparsity and structural gaps inherent in the original dataset but also substantially extends its complexity frontier—surpassing the original in both scale and quality. The broader coverage and increased diversity enabled by data augmentation are reflected in Figure 4. Using a 2D t-SNE projection, the figure illustrates both database-level and feature-level distributions. The left panel indicates that auxiliary databases expand the coverage beyond the original dataset, while the right panel reveals a more diverse spread in feature representations derived from the statistics in Table I.
III-C Database Interaction Implementation
Multiple stages of the Text2SQL-Flow data augmentation pipeline require frequent interactions with the underlying database, including executing SQL queries for validity checking and accessing schema information for query generation. To support multiple database systems and improve implementation robustness, we implement a database interaction layer that abstracts database-specific driver details and provides a unified interface to upper-layer components. This interface supports batched SQL execution and schema access through standardized APIs, enabling seamless integration of different relational database systems (e.g., SQLite, MySQL) without modifying the core pipeline logic. For efficiency, schema information is cached after the first access to avoid redundant metadata queries during large-scale SQL generation.
III-D Synthetic Data for Open-Source Models
The pairs $\langle p_{i},cot_{i}\rangle$ generated above serve as high-quality training data $\mathcal{D}$. During SFT, the model is trained on the full output sequence using standard causal language modeling:
@@ III-D Synthetic Data for Open-Source Models
where the input $x$ corresponds to the prompt $p_{i}$, and the output $y$ represents the complete CoT sequence $cot_{i}$. The model parameters are denoted by $\theta$. By optimizing this loss over the entire output sequence, the model is encouraged to learn interpretable reasoning pathways and improve its performance on the Text-to-SQL task.
III-E Synthetic Data for Prompting-based Methods
@@
However, existing methods for retrieving few-shot examples fail to provide effective guidance because similarity has not been sufficiently exploited. In this case, we propose a structure-aware few-shot retrieval method for Text-to-SQL. The core idea is to map both NL questions and SQL queries into a unified, detail-abstracted semantic space, where retrieval is performed based on structural similarity. Specifically, inspired by the masking strategy in DAIL–SQL [10], we apply standardized masking to schema-dependent elements (e.g., table/column names), string literals, and numeric constants in both questions and SQL queries, denoted as $\text{Mask}(q)$ and $\text{Mask}(\text{s})$, respectively. Building upon this, we develop a structure-aware embedding model to compute semantic similarity between masked NL questions and masked SQL queries. Given a target question $q_{t}$, we retrieve the top-$k$ examples from a retrieval pool $B=\{(q_{i},s_{i})\}_{i=1}^{N}$, which we refer to as a knowledge base. $B$ consists of NL–SQL pairs from the training dataset, with no overlap with the test dataset. $B$ is stored as an unordered collection of independent data pairs, rather than a structured knowledge graph.
@@
In our implementation, we adopt the Qwen3-Embedding architecture [55], employing a causal-attention-based LLM as the embedding encoder. Both the knowledge base and the training data for the embedding model are derived from the SQLFlow dataset we constructed. The input format concatenates the task instruction with the masked question as follows: $\texttt{\{Instruction\} \{Mask}(q_{i})\texttt{\}}$ followed by an [EOS] token. The final sentence embedding is obtained from the hidden state of the [EOS] token in the last layer of the encoder. Our method retrieves few-shot examples that are highly aligned with the target query in SQL structure. This enhances the few-shot reasoning capability of closed-source LLMs on the Text-to-SQL task.